# Beyond Prompt Content: Enhancing LLM Performance via Content-Format Integrated Prompt Optimization

### v2

@@ 1 Introduction
Figure 1: The crucial role of prompt formatting and its interaction with content. (A): Model-specific format biases: Illustrates the performance sensitivity of two LLMs to different format styles on the GSM8K task, showing substantial variability in the effectiveness of 10 randomly selected formats. (B): For seven different prompt contents evaluated across 24 distinct formats, performance variations show the complex, interdependent relationship between prompt content and structure, demonstrating that no single format universally maximizes effectiveness.
@@
However, manual design of prompts presents significant challenges, primarily due to the high sensitivity of LLMs to subtle variations in prompt characteristics, including both textual content and structural format (Jiang et al., 2022; Zamfirescu-Pereira et al., 2023; Salinas and Morstatter, 2024). These sensitivities are further complicated by variations across different models and tasks Zhuo et al. (2024); Sclar et al. (2024). To alleviate these difficulties, automated prompt optimization techniques, often leveraging the power of LLMs themselves, have proven to be an effective approach to adapt and refine prompts Pryzant et al. (2023); Schnabel and Neville (2024); Yang et al. (2024). However, existing research primarily focuses on optimizing prompt content, while overlooking a critical and largely unexplored dimension: the prompt formatting.
Our preliminary investigations, as illustrated in Figure 1, provide valuable insights into the role of prompt format in prompt optimization. We have observed that different LLMs display distinct preferences, with some formats performing well on one model but failing on another. This suggests sophisticated, model-specific format biases Sclar et al. (2024). Furthermore, we have identified a complex interplay between prompt content and format, where no single format consistently outperforms others across all contents. This lack of a universally optimal format highlights the impracticality of predefining format, and underscores the need for a joint optimization approach that treats prompt content and format as interdependent variables.
@@
To address these limitations, we introduce Content-Format Integrated Prompt Optimization (CFPO), an innovative methodology that concurrently optimizes both prompt content and format through an iterative refinement process. CFPO employs distinct optimization strategies tailored to the unique search spaces of content and format. Content optimization is guided by performance feedback and Monte Carlo sampling, leveraging natural language mutations to enhance prompt effectiveness. For format optimization, CFPO explores a discrete set of format options through a dynamic exploration strategy designed to identify optimal formats without requiring a prior knowledge.
Specifically, CFPO’s format optimizer leverages the principles of structured thinking, operating along two key dimensions: the Prompt Renderer, which governs the organizational structure of all components within a prompt (He et al., 2024), and the Query Format, which dictates the presentation of in-context learning examples and queries (Voronov et al., 2024a; Salinas and Morstatter, 2024). By integrating these two dimensions, CFPO defines a structured template that effectively distinguishes between content and format types, enabling the efficient identification of high-performing prompts.
Our primary contributions are threefold: (1) We propose CFPO, an innovative approach to simultaneously optimizes prompt content and format using an iterative process. (2) We introduce an efficient strategy for dynamic format optimization that generates new formats in an iterative manner and evaluates formats instance through a scoring system to select the best option. (3) Through extensive evaluations across diverse tasks and multiple open-source LLMs, we demonstrate that CFPO consistently improves LLM performance in a measurable and effective manner.
2 Related Work
Optimization via LLM The remarkable capacity of LLMs has been demonstrated in various tasks as optimizers, leveraging their ability to enhance performance, such as code generation (Haluptzok et al., 2023; Zelikman et al., 2024; Askari et al., 2024), tool-making (Cai et al., 2024), and agent system design (Hu et al., 2024). However, recent studies indicate that LLMs face significant challenges in achieving completely automatic optimization. These models often rely on human intervention for designing workflows and struggle with tasks requiring complex decomposition and iterative refinement (Zhang et al., 2024; Li et al., 2024).
Automatic Prompt Optimization Automatic prompt optimization plays a crucial role in enhancing the performance of LLMs by refining prompts without requiring human intervention. Various approaches have been explored to search for the optimal prompt, including reinforcement learning (Zhang et al., 2023), Monte Carlo Search (Zhou et al., 2023), Monte Carlo Tree Search (MCTS) (Wang et al., 2024b), feedback-based methods (Pryzant et al., 2023; Das et al., 2024), and agent-driven frameworks (Wang et al., 2024a; Khattab et al., 2024; WHO, 2023). While these methods focus on optimizing the overall prompt, they often lack the capability for fine-grained modifications. (Khattab et al., 2024; Schnabel and Neville, 2024) introduce phrase-level mutations, but they fail to address format mutations or implement them in a systematic manner.
Prompt structure and format Structured prompting, which organizes prompts into distinct components such as instructions, examples, and queries, holds significant potential in prompt engineering (Fernando et al., 2023). Empirical rules for prompt design always lack integration with automatic optimization techniques, limiting their scalability and effectiveness (Nigh, 2023; Google, 2024). Frameworks like LangGPT (Wang et al., 2024a) have introduced structured prompting paradigms, emphasizing reusable designs inspired by programming principles. However, these efforts primarily focus on content-level refinements and fail to adequately address the critical role of prompt formatting. Studies have highlighted the impact of formatting on prompt performance (Salinas and Morstatter, 2024). Sclar et al. (2024) revealed that modifications to separators and spacing within a query could substantially impact performance. He et al. (2024) reveals that the format of prompts significantly impacts GPT-based models’ performance, with no single format excelling universally. Voronov et al. (2024b) focus on the format of few-shot examples and suggests that it is beneficial to maintain a consistent format across examples. However, despite the recognition of formatting’s importance, there remains a lack of comprehensive understanding regarding the optimization of prompt format in a systematic manner.
3 CFPO: Content-Format Integrated Prompt Optimization
As we have established, the effectiveness of LLMs is profoundly influenced by both the content and format of prompts. Existing automated prompt optimization methods have largely overlooked the format dimension, which exhibits a strong model bias. To address this critical limitation, we introduce Content-Format Integrated Prompt Optimization (CFPO) framework that jointly optimizes both prompt content and format. This contrasts with prior approaches focusing solely on content optimization. Our goal is to identify an optimal prompt $p^{*}$, comprising both content ($c^{*}$) and format ($f^{*}$), that maximizes performance on an evaluation dataset $\mathcal{D}$, given by:
$p^{*}:(c^{*},f^{*})=\arg\max_{c\in\mathcal{L},f\in\mathcal{F}}m(c,f|\mathcal{D}),$ (1)
within the coherent natural language space $\mathcal{L}$ and the space of all possible formats $\mathcal{F}$, guided by a metric function $m(\cdot)$ that assesses the prompt’s quality.
To effectively search this complex space, CFPO employs a two-pronged iterative approach, detailed in Figure 2, that consists of two concurrently-run optimizers: a Component-wise Content Optimizer and a Format Optimizer. The content optimizer refines the textual content of a prompt, while the format optimizer explores the structural arrangement of its elements. Importantly, our framework acknowledges the inherent interdependence of content and format and thus iterates between optimizing them, to find their optimal combination. The following sections detail our structured prompt template (Section 3.1), our innovative format optimization approach (Section 3.2), and finally our integrated optimization process (Section 3.3).
@@ 3.1 Structured Prompt Template
To enable fine-grained and targeted optimization, our framework adopts a structured prompt template inspired by guidelines from  OpenAI (2024b) and Google (2024). This template decomposes prompts into distinct functional components, facilitating both analysis and selective mutations. Specifically, our template divides a prompt into content-based components and format-based components, as illustrated in Figure 3.
The Content-based Components are:
@@
Few-shot Examples provide contextual learning patterns, consisting of:
  • Examples: specific instances pertinent to the task, including inputs and expected outputs.
  • Example Hinter (optional): a brief hint indicating that examples segment will follow, e.g., ’Here are some examples:’.
  • CoT Hinter (optional): encourages a chain-of-thought reasoning process, e.g., ’Let’s think step by step’.
@@
The Format-based Components are:
@@
The formulation of the structured prompt template is fundamental to our optimization approach. This design yields two key advantages: first, it facilitates a structured component functionality where each part serves a specific purpose, promoting a more organized prompting framework; second, it enables fine-grained optimization by decoupling format from content, thus allowing targeted and precise modifications of individual components.
3.2 Format Optimizer Design
The key aspect of our work is the format optimization methodology. To efficiently explore the extensive range of prompt formats, the CFPO format optimizer adopts an approach that utilizes a format pool with a scoring system and an LLM-assisted format generation module. It strategically explores, evaluates, and refines formatting choices, all while learning from previous iterations.
@@ 3.2.1 Format Pool with Scoring System
The format pool is designed to hold the format configurations we use to generate prompts. As shown in Figure 4, these configurations are separated into two dimensions: the Prompt Renderer, which dictates the overall structure of the prompt, and the Query Format, which governs the rendering of in-context examples and queries. This distinction allows us to explore both macro and micro-level formatting variations.
To dynamically evaluate the potential of each format, we developed a scoring system for assessing the performance of each format $f$, represented as $Q(f)$. This system updates the performance score of $f$ across various prompt contents using the formula $Q(f)\leftarrow Q(f)+\sum_{c}m(c,f)$, where $c$ represents each content instance in current round. Additionally, we maintain $N(f)$ to count the number of times a format has been visited, which facilitates score normalization.
To initialize the exploration, we constructed a initial format search space $\mathcal{F}$, comprising a set of predefined commonly used formats, as illustrated in Figure 4. We also incorporate diverse variations of these predefined formats into the initial search space, such as adjustments to spacing, punctuation, and the use of special symbols. This establishes a starting point for our optimization.
@@ 3.2.2 LLM-assisted Format Generation
The variability of format space requires an automated process for effective expansion and exploration. To that end, we introduce an LLM-based format generator, $LLM_{f\_gen}$, which autonomously generates new formats based on information in the existing format pool.
This evolutionary approach integrates the format generation into each optimization round, allowing for the creation of new and potentially beneficial formats. To enhance the efficiency of this process, we guide the LLM towards more promising areas by informing it of the performance function, $\frac{Q(f)}{N(f)}$. This iterative process not only diversifies the format pool but also ensures that our system can adapt to and incorporate a wide range of formats, thereby enhancing its utility and effectiveness. More detailed information of our format generation process is provided in the Appendix A.2.
@@ 3.2.3 Search Format via Format Optimizer
$UCT(f)=\frac{Q(f)}{N(f)}+\alpha\sqrt{\frac{\sum_{f}N(f)}{N(f)}}$ (2)
The overall process, outlined in Algorithm 1, selects $2k$ formats for evaluation in each optimization round: $k$ promising formats from the pool (based on UCT score), and $k$ new formats generated by the $LLM_{f\_gen}$. The selected formats from both the existing pool ($\mathcal{F}_{select}$) and the new generated pool ($\mathcal{F}_{gen}$) are then evaluated using a predefined metric function $m(\cdot)$, and the best-performing format among the tested candidates will be identified. The result is then incorporated into the pool for future iterations.
@@
By iteratively evaluating formats, the format optimizer ensures a balance between exploring new formats and refining current ones, converging to the best format configuration.
@@ 3.3 Integrated Optimizer Design
CFPO orchestrates the Component-wise Content Optimization and Format Optimization within an iterative framework to jointly optimize content and format. This iterative process (illustrated in Figure 2) is key to our methodology.
Component-wise Content Optimization: This stage employs two primary strategies for mutating the content of prompts. The first strategy is case-diagnosis and revision, leveraging test cases to assess the efficacy of the current prompt. The outcomes of these test cases, including both correct and incorrect samples, are analyzed by the LLM optimizer. This optimizer evaluates the performance and pinpoints specific components in need of optimization. Subsequently, targeted feedback is applied to these identified components for enhancement, resulting in improved prompts. For example, if the output is not in the specified format, the output format component will be altered. Additionally, a Mote-Carlo sampling strategy is employed to enhance the optimization robustness by generating synthetic content with same semantics for randomly selected components. After this step, we select top-performing content candidates based on an evaluation dataset for the next stage.
Format Optimization: As discussed in Section 3.2, this stage identifies the most effective format for each candidate prompt content from the previous content-optimization step. The format optimizer applies our dynamic format exploration and evaluation process, tracking performance, and updating the format pool’s scoring system. The Format Optimizer meticulously tracks the performance of all evaluated formats, providing valuable insights to guide the selection of formats in subsequent iterations. Simultaneously, it retains only the most effective format for each prompt, ensuring the diversity of prompt content candidates during beam search.
In summary, the two optimizers work in tandem, leveraging the strengths of the LLM to facilitate swift adaptation and customization. Importantly, this iterative process allows for the optimization of format and content, thereby significantly enhancing the quality of the generated prompts.
@@
ARC-Challenge
Baseline 67.15 73.81 75.94 84.39
GRIPS 77.05 77.90 79.61 87.46
APE 75.85 77.05 78.67 87.63
ProTeGi 76.54 77.22 79.86 87.54
SAMMO 77.22 77.13 79.86 87.03
CFPO (Ours) 79.35 78.50 80.63 88.23
Big-Bench Classification
Baseline 56.00 64.00 70.00 54.00
GRIPS 86.00 67.00 84.00 69.00
APE 73.00 65.00 60.00 63.00
ProTeGi 83.00 81.00 82.00 76.00
SAMMO 86.00 80.00 86.00 78.00
CFPO (Ours) 94.00 90.00 91.00 87.00
Table 1: Main results on math reasoning tasks and commonsense reasoning tasks.

### v3

@@ 1 Introduction
Figure 1: The crucial role of prompt formatting and its interaction with content. (A): Model-specific format sensitivity: LLM performance varies significantly across different prompt formats on GSM8K task. (B): Content-format interdependence: The optimal format for a prompt depends on its content, highlighting the need for joint optimization.
Despite its significance, manual prompt design is fraught with challenges. LLMs are notoriously sensitive to subtle variations in prompt phrasing and structure, with performance often differing markedly across models and tasks based on these nuances (Zhuo et al., 2024a; Chatterjee et al., 2024; Jiang et al., 2022; Salinas and Morstatter, 2024; Zhuo et al., 2024b; Sclar et al., 2024). To alleviate these difficulties, automated prompt optimization techniques, often leveraging enhanced LLMs to optimize prompts, have proven to be effective to adapt and refine prompts Pryzant et al. (2023); Schnabel and Neville (2024); Yang et al. (2024).
However, existing research primarily focuses on optimizing prompt content, while overlooking a critical and unexplored dimension: the prompt formatting. Prompt formatting refers to the arrangement and presentation of prompt content while preserving its semantic meaning. As LLMs are applied to increasingly complex tasks, structuring prompts into distinct components (e.g., instructions, examples, queries) becomes paramount for effectively conveying the desired task and context. Thus, the manner in which a prompt is formatted can significantly impact performance.
Our preliminary investigations (Figure 1) highlight the significant impact of prompt format on LLM performance. We observed that different LLMs exhibit distinct format preferences, with formats performing well on one model sometimes failing on another. This underscores the existence of sophisticated, model-specific format biases Sclar et al. (2024). Furthermore, even within a single LLM, the optimal format varies depending on the specific prompt content. This complex interplay between content and format suggests that a one-size-fits-all approach to prompt formatting is unlikely to succeed, highlighting the need for joint optimization strategies that consider content and format as interdependent variables.
To address these limitations, we introduce Content-Format Integrated Prompt Optimization (CFPO), an innovative methodology that concurrently optimizes both prompt content and format through an iterative refinement process. CFPO employs distinct optimization strategies tailored to the unique search spaces of content and format. Content optimization is guided by performance feedback and Monte Carlo sampling, leveraging natural language mutations to enhance prompt effectiveness. For format optimization, CFPO explores a discrete set of format options through a dynamic exploration strategy designed to identify optimal formats without requiring prior knowledge.
Specifically, CFPO’s format optimizer draws upon principles of structured thinking and defines a hierarchical template that clearly demarcates content elements from their formatting attributes. This allows for targeted optimization of both intra-component styling (e.g., how an example is presented) (Voronov et al., 2024a; Salinas and Morstatter, 2024) and inter-component arrangement (e.g., the order and connectors between components) (He et al., 2024), adapting to the specific needs of different prompt components and their interactions.
Our primary contributions are threefold:
  • We propose CFPO, an innovative approach to simultaneously optimizes prompt content and format using an iterative process.
  • We introduce an efficient strategy for dynamic format optimization that iteratively generates and evaluates format candidates using a scoring system to select superior options.
  • We conduct extensive evaluations across diverse tasks and open-source LLMs, showing that CFPO consistently enhances model performance in a measurable and effective way.
2 Related Work
@@
2.1 Optimization via LLM
The remarkable capacity of LLMs has been demonstrated in various tasks as optimizers, leveraging their ability to enhance performance, such as code generation (Haluptzok et al., 2023; Zelikman et al., 2024; Askari et al., 2024), tool-making (Cai et al., 2024), and agent system design (Hu et al., 2024). However, recent studies indicate that LLMs face significant challenges in achieving completely automatic optimization. These models often rely on human intervention for designing workflows and struggle with tasks requiring complex decomposition and iterative refinement (Zhang et al., 2024; Li et al., 2024).
2.2 Prompt Optimization Techniques
Automatic Prompt Optimization Automatic prompt optimization plays a crucial role in enhancing the performance of LLMs by refining prompts without requiring human intervention. Various approaches have been explored to search for the optimal prompt, including reinforcement learning (Zhang et al., 2023), Monte Carlo Search (Zhou et al., 2023), Monte Carlo Tree Search (MCTS) (Wang et al., 2024b), and agent-driven frameworks (Wang et al., 2024a; Khattab et al., 2024; WHO, 2023). Notably, feedback-based methods have emerged as a significant approach (Pryzant et al., 2023; Das et al., 2024). These methods iteratively evaluate prompts on a batch of evaluation data, using error cases to guide subsequent mutation and improvement (Pryzant et al., 2023; Das et al., 2024). In each iteration, the prompt optimizer evaluates the prompt on the evaluation set and feeds the incorrectly predicted test cases back to the optimizer to guide improvement in the next mutation step. However, existing automatic prompt optimization techniques often lack the capacity for fine-grained modifications. While (Khattab et al., 2024; Schnabel and Neville, 2024) introduce phrase-level mutations, a systematic approach to optimizing prompt format is lacking. This leaves a gap to comprehensively adapt prompt structure for optimal performance.
Prefix Tunning offers an effective strategy for adapting large language models (LLMs) to specific tasks by learning continuous, task-specific vectors that are prepended to the input sequence Li and Liang (2021); Wang et al. (2023); Guo et al. (2024); Gu et al. (2022); Wang et al. (2023). While demonstrating impressive performance gains, a key limitation of Prefix Tuning lies in its reliance on access to the LLM’s internal parameters for training. This requirement presents a significant obstacle when working with black-box LLMs accessed through APIs or in resource-constrained environments where full fine-tuning is infeasible.
2.3 Prompt Structure and Sensitivity
Structured Prompting, which organizes prompts into components like instructions and examples, is a widely recommended practice (OpenAI, 2024b; Google, 2024). Frameworks like LangGPT (Wang et al., 2024a) further advance this paradigm by introducing reusable prompt designs inspired by software engineering principles, showcasing the potential of structured approaches (Fernando et al., 2023). While these efforts have primarily focused on optimizing the content and organization of prompt components, less attention has been paid to the impact of prompt formatting within this structured context.
Sensitivity of Prompt Variations LLMs exhibit significant sensitivity to prompt variations. Studies have shown that even semantically similar but unseen instructions can lead to performance degradation (Sun et al., 2023; Mizrahi et al., 2024). To address this brittleness, researchers have proposed methods and metrics to systematically evaluate and understand prompt sensitivity (Zhuo et al., 2024b; Chatterjee et al., 2024). These findings underscore the necessity of prompt optimization for robust LLM performance.
Format Sensitivity of Prompt Studies have highlighted the impact of formatting on prompt performance (Salinas and Morstatter, 2024). Sclar et al. (2024) revealed that modifications to separators and spacing within a query could substantially impact performance. He et al. (2024) reveals that the format of prompts significantly impacts GPT-based models’ performance, with no single format excelling universally. Voronov et al. (2024b) focuses on the format of few-shot examples and suggests that it is beneficial to maintain a consistent format across examples. Despite the growing recognition of formatting’s influence, current prompt engineering practices rely heavily on empirical observations and lack systematic design principles.
3 CFPO: Content-Format Integrated Prompt Optimization
The effectiveness of LLMs is profoundly influenced by both the content and format of prompts. To address this, we propose Content-Format Integrated Prompt Optimization (CFPO), a novel framework designed to jointly optimize prompt content and structure for enhanced LLM performance. CFPO, illustrated in Figure 2, employs a dual-optimizer approach in each iteration. The subsequent sections provide a detailed explanation of our methodology, beginning with the structured prompt template we utilize (Section 3.1), followed by a description of the Component-wise Content Optimization (Section 3.2), the Format Optimization (Section 3.3), and finally, the integration of these optimizers into the complete CFPO framework (Section 3.4). Implementation details and meta-prompts are provided in Appendix A.
@@ 3.1 Structured Prompt Template
To facilitate fine-grained and targeted optimization, our framework adopts a structured prompt template inspired by prompt engineering guidelines from  OpenAI (2024b) and Google (2024). Our template decomposes prompts into distinct functional components, allowing for both detailed analysis and selective modification. The template distinguishes between content-based and format-based components, providing flexibility and adaptability to diverse user needs. A set of common and generally applicable elements is illustrated in Figure 3.
The Content-based Components define the information provided to the LLM, including:
@@
  • Few-shot Examples provide specific instances for contextual learning patterns.
@@
The Format-based Components dictate how the prompt is assembled and presented, including:
@@
Note that this template is highly adaptable. Users can readily adjust the template by adding or deleting components, or incorporating additional formatting elements (e.g., tables, structured documents) within existing components to suit their specific requirements. By decoupling format from content, this structured template empowers users to perform targeted and precise optimizations, leading to improved prompt effectiveness.
3.2 Component-wise Content Optimization
Building upon existing prompt optimization strategies, which often rely on either feedback-driven refinement or Monte-Carlo sampling (Section 2.2), CFPO introduces a more targeted and efficient approach to content optimization. While feedback-driven methods diagnose weaknesses through failure analysis, and Monte-Carlo sampling diversifies perspectives with semantic variations, CFPO innovates along two key dimensions.
First, CFPO expands the diagnostic phase beyond traditional failure analysis by incorporating correct cases. This allows the LLM to identify and reinforce successful aspects of the prompt, complementing error correction. Second, CFPO adopts a component-wise optimization strategy. Instead of treating the prompt as a monolithic block, CFPO targets specific content-based components for individual refinement, enabling more precise and effective optimization.
3.3 Format Optimization Design
To efficiently navigate the prompt format space, CFPO incorporates a format optimizer based on a dynamic format pool and an LLM-assisted generation module. The optimizer iteratively explores, evaluates, and refines formatting choices.
@@ 3.3.1 Format Pool with Scoring System
The format pool maintains a diverse collection of prompt format configurations, organized hierarchically to represent variations at both macro (e.g., overall prompt structure) and micro (e.g., component-internal rendering) levels. In this work, we prototype this structure with two distinct renderers: a Prompt Renderer defining the global structure, and a Query Format governing the rendering of individual components, such as examples and queries, as shown in Figure 4.
Each format $f$ in the pool is associated with a performance score, $Q(f)$, which cumulatively reflects its effectiveness across different prompt contents. When a format $f$ is evaluated on a new content $c$, its score is updated as follows: $Q(f)\leftarrow Q(f)+m(c,f)$, where $m(c,f)$ is a task-specific metric function. For example, in reasoning tasks, $m(\cdot)$ represents the average accuracy (0 or 1) on an evaluation dataset. The number of evaluations for each format, $N(f)$, is also tracked to enable score normalization and fair comparison.
To initiate the format exploration process, we define an initial format search space, $\mathcal{F}$, consisting of a set of predefined and commonly used formats (Figure 4), along with diverse variations that introduce subtle changes (e.g., spacing, punctuation, special symbols). This initial format pool serves as the foundation for subsequent exploration.
@@ 3.3.2 LLM-assisted Format Generation
To overcome the limitations of a static format pool and promote continuous adaptation, we introduce an LLM-based format generator, $LLM_{f\_gen}$. This module autonomously generates novel formats by leveraging information about the existing formats in the pool and their performance characteristics. By dynamically creating new formats, we aim to escape local optima and discover potentially superior formatting strategies.
In each optimization round, the format generator is provided with the existing formats and their normalized performance scores, $\frac{Q(f)}{N(f)}$, and tasked with creating new formats or variations that might improve performance. This iterative process not only diversifies the format pool but also ensures that our system can adapt to and incorporate a wide range of formats, thereby enhancing its utility and effectiveness. Additional details and generated examples about the format generation process are available in Appendix A.3 and Appendix D.1.
@@ 3.3.3 Search Format via Format Optimizer
$UCT(f)=\frac{Q(f)}{N(f)}+\alpha\sqrt{\frac{\sum_{f}N(f)}{N(f)}}$ (1)
The overall process, outlined in Algorithm 1, selects $2k$ formats for evaluation in each optimization round: $k$ existing formats from the pool (based on UCT score), and $k$ new formats generated by the $LLM_{f\_gen}$. The selected formats from both the existing pool ($\mathcal{F}_{select}$) and the newly generated pool ($\mathcal{F}_{gen}$) are then evaluated using a predefined metric function $m(\cdot)$, and the best-performing format among the tested candidates will be identified. The result is then incorporated into the pool for future iterations. By iteratively evaluating formats, the format optimizer ensures a balance between exploring new formats and refining current ones, converging to the best format configuration.
@@ 3.4 Integrated Optimizer Design
CFPO integrates content and format optimization within an iterative loop, as illustrated in Figure 2. In each iteration, the content optimizer proposes candidate prompts, and the format optimizer identifies the most effective format for each candidate. This ensures that each prompt benefits from the most effective formatting. Furthermore, effective formats can improve the diversity and performance of content candidates, thereby helping content optimizer for beam search in the next iteration.
In summary, the content and format optimizers work in tandem, leveraging the LLM to enable rapid adaptation and customization. This iterative and collaborative process optimizes both the content and the format of prompts, leading to significant improvements in prompt quality.
@@
Big-Bench Classification
Baseline 56.00 64.00 70.00 54.00
GRIPS 86.00 67.00 84.00 69.00
APE 73.00 65.00 60.00 63.00
ProTeGi 83.00 81.00 82.00 76.00
SAMMO 86.00 80.00 86.00 78.00
CFPO (Ours) 94.00 90.00 91.00 87.00
ARC-Challenge
Baseline 67.15 73.81 75.94 84.39
GRIPS 77.05 77.90 79.61 87.46
APE 75.85 77.05 78.67 87.63
ProTeGi 76.54 77.22 79.86 87.54
SAMMO 77.22 77.13 79.86 87.03
CFPO (Ours) 79.35 78.50 80.63 88.23
@@
Table 1: Main results of CFPO and state-of-the-art methods on four datasets.